OpenStudy (anonymous):

A city has been declining in population for the last 50 years. In the year 2000, the population was 571 thousand, and was declining at a rate of 0.77% per year. If this trend continues: (a) Give a formula for the population of this city, P, in thousands, as a function of t, the number of years since 2000. P= (b) What is the predicted population in 2010? The predicted population is thousand people. (c) To two decimal places, estimate t when the population is 450 thousand people. When there are 450 thousand people in this city, t≈

OpenStudy (whpalmer4):

what have you done so far?

OpenStudy (anonymous):

Well, actually,P(t)=571000(.23)^t

OpenStudy (anonymous):

\[\large 450=571e ^{0.0077t}\]

OpenStudy (anonymous):

Solve for t.

OpenStudy (anonymous):

I haven't learned about the #e yet.

OpenStudy (anonymous):

Exponential is e.

OpenStudy (anonymous):

Alright, well thank you for trying to help me! I am just not allowed to use e in my equation because that is the wrong answer according to my teacher.

OpenStudy (wolf1728):

You don't necessarily have to use the constant 'e'. Population = 571,000 * .9923^n where n= (present year minus 2000) So, for an example, the population in 2003 would be: 571,000 * .9923^3 which equals 557,911 They want the formula to provide answers in thousands: a) So we have the formula: Population (in thousands) = (571 * .9923^t) where 't' equals the number of years since 2000 b) The population in 2010 is Population (in thousands) = (571 * .9923^10) Population (in thousands) = 528.525592559 Population (in thousands) (rounded) = 528.526 c) To find when population will be 450,000 we have to manipulate the formula: 450 = (571 * .9923^x) (450 / 571) = .9923^x To find the time (or 'x') we must take logarithms of both sides: log (450 / 571) = x * log (.9923) log (0.7880910683) = x * log (.9923) -0.1034235945 = x * -0.0033570086 x = -0.1034235945 / -0.0033570086 x = 30.8082597997 years or rounding this to 2 decimal places x = 30.81 years and see? It's just that simple - LOL

OpenStudy (anonymous):

Haha. Thank you SO much! You are a lifesaver!

OpenStudy (whpalmer4):

"wolframalpha" didn't get the formula quite right, so it's just as well that you couldn't use it :-) If you set up the decay with the exponential function, the value of \(\lambda\) should be negative in order to get a positive value of \(t\). \[450 = 591 e^{-0.0077298t}\]is what you would solve for \(t\). \[\ln(\frac{450}{571}) = -0.0077298t\]\[-0.238142 = -0.0077298t\]\[t = 30.8083\]